Suppose the frequency of a system is measured to be 2rad/s and after the addition of of another mass 1kg, the frequency is measured to be 1rad/s. Calculate the mass and stiffness.


1
Expert's answer
2021-04-21T07:05:40-0400

w = (k/m)1/2

2 rad/s = (k/x)1/2

4 = k/x

1 rad/s = (k/(x+1))1/2

1 = k/(x+1)

k = 4x

k = x+1

4x = x +

x = 1/3 kg

k = 4 x 1/3 = 1.33 N/m